Suppose that the distance a car travels varies directly with the amount of gasoline it uses. A certain car uses 9 gallons of gasoline to travel 234 miles. How many miles can the car travel if it has 23 gallons of gasoline?
Found 2 solutions by vleith, stanbon:
Answer by vleith(2825)
Answer by stanbon(57424)

Suppose that the distance a car travels varies directly with the amount of gasoline it uses. A certain car uses 9 gallons of gasoline to travel 234 miles. How many miles can the car travel if it has 23 gallons of gasoline.
----
d = kg
----
A certain car uses 9 gallons of gasoline to travel 234 miles.
---
234 = k*9
So, k = 234/9
----
Equation for that car
d = (234/9)g
------
How many miles can the car travel if it has 23 gallons of gasoline.
------
d = (234/9)*23
d = 598 miles
===================
Cheers,
Stan H.